A projectile on Earth is fired straight upward so that its distance (in feet) above the ground t seconds after firing is given by s (t) = 16tē + 400t
Find the maximum height it reaches and the number of seconds it takes to reach that height.

Answer by ewatrrr(24785)


Hi
s (t) = -16t^ + 400t |s' = -32t + 400, t = 12.5
= -16(t - 400/32)^2) + 2500 |Completing Square
= -16(t - 12.5)^2) + 2500
Hmax = 2500ft in 12.5sec